OpenStudy (lgbasallote):

an adventurous archaelogist (m=85.0 kg) tries to cross a river by swinging from a vine. the vine is 10.0 m long, and his speed at the bottom of the swing is 8.00 m/s. the archaelogist doesnt know that the vine has a breaking strength of 1,000 N. does he make it across the river without falling in?

mathslover (mathslover):

lgba i want u to be here for a while please

mathslover (mathslover):

@lgbasallote

OpenStudy (lgbasallote):

yes im here

mathslover (mathslover):

ok ! so .. can u calculate the weight of adventurus archaelogist ?

mathslover (mathslover):

i am taking adventurous archaelogist as X ... ok ?

OpenStudy (anonymous):

w = mg

mathslover (mathslover):

Weight = mg .. Weight = 85 kg. * 9.8 \(ms^{-2}\)

OpenStudy (lgbasallote):

ys i got the weight

OpenStudy (lgbasallote):

i think it's 522 if i remember right

mathslover (mathslover):

hence we have weight = 833 N . good .. Now do you the equation of centripetal acceleration ?

OpenStudy (anonymous):

a = v^2/r

mathslover (mathslover):

85 * 9.8 = 833 N ............................. not 522 ... check it again @lgbasallote

mathslover (mathslover):

saurav shakya ... Please let lgba .. do ..

OpenStudy (lgbasallote):

oops lol 522 was the one i did for centripetal force

OpenStudy (lgbasallote):

83.3 N

mathslover (mathslover):

833 N now centripetal acceleration = \(\frac{v^2}{R}\)

mathslover (mathslover):

right ?

OpenStudy (lgbasallote):

yup

mathslover (mathslover):

ok so ... R = 10 m ... and v = 8 \(ms^{-1}\)

OpenStudy (anonymous):

now, tension = weight of the adventurous archaelogist + centripetal force

OpenStudy (lgbasallote):

sounds legit

mathslover (mathslover):

centripetal acceleration = \(\frac{64}{10}=6.4\)

mathslover (mathslover):

centripetal acceleration = 6.4 \(ms^{-2}\)

OpenStudy (lgbasallote):

uh huh

mathslover (mathslover):

sum of the forces on X = mass * acceleration = 85 kg * 6.4 \(ms^{-2}\) =544 N

OpenStudy (anonymous):

is it so

OpenStudy (lgbasallote):

it's not 522?

OpenStudy (lgbasallote):

so i add those forces?

OpenStudy (lgbasallote):

gravity and centripetal?

mathslover (mathslover):

T - mg = ma = 544 N mg = 833 N T = 833 N + 544 N

OpenStudy (lgbasallote):

dont ou mean 83.3?

mathslover (mathslover):

yes by adding them we get T .. that is above 1000 N Hence ..... the vine will break ...

OpenStudy (anonymous):

so he will not make it

OpenStudy (lgbasallote):

uhmm oops i made m = 8.5

OpenStudy (lgbasallote):

okay i got it

OpenStudy (lgbasallote):

thanks

mathslover (mathslover):

nice to hear. .. best of luck dude